(1 point) An investor purchases two bonds with the following properties: Bond 1: Has a face value $1000 and is redeemable at par. Pays coupons annually at a rate of 8% annual and was purchased for $1084.58. Bond 2: Has a face value $1000 and is redeemable at par. Pays coupons annually at a rate of 6.4% annual and was purchased for $996.7. If both bonds mature in the same number of years and the investor yields the same rate on both bonds, find the yield rate as a percent correct to 2 decimal places). Answer:
Answer #1

Yield To Maturity(YTM) = (interest per period+ ((Redemption price - Current market price) / life remaining to maturity)) / ((.4*Redemption price)+ (.6*Current market price))

Let n be the life remaining to maturity.

YTM of Bond 1 = YTM of Bond 2

((1000*8%)+((1000-1084.58)/n)) / (.4*1000+.6*1084.58) = ((1000*6.4%)+((1000-996.7)/n)) / (.4*1000+.6*996.7)

(80+((-84.58)/n)) / 1050.748 = (64)+((3.3)/n)) / 998.02

(80+((-84.58)/n)) * 998.02 = (64)+((3.3)/n)) * 1050.748

79841.6-(84412.5316/n) = 67247.872+(3467.4684/n)

3467.4684/n + 84412.5316/n = 79841.6 - 67247.872

87880/n = 12593.728

n = 87880/12593.728

= 6.97807670612

life remaining to maturity = 7

Put the value of n to any of the equation

YTM = (80+((-84.58)/7)) / 1050.748

= (80-12.0828571429) / 1050.748

= 67.9171428571/1050.748

= 0.06463694706

= 6.46%